### v2

# A Fast and Optimal Learning-based Path Planning Method for Planetary Rovers

Abstract
Intelligent autonomous path planning is crucial to improve the exploration efficiency of planetary rovers. In this paper, we propose a learning-based method to quickly search for optimal paths in an elevation map, which is called NNPP. The NNPP model learns semantic information about start and goal locations, as well as map representations, from numerous pre-annotated optimal path demonstrations, and produces a probabilistic distribution over each pixel representing the likelihood of it belonging to an optimal path on the map. More specifically, the paper computes the traversal cost for each grid cell from the slope, roughness and elevation difference obtained from the DEM. Subsequently, the start and goal locations are encoded using a Gaussian distribution and different location encoding parameters are analyzed for their effect on model performance. After training, the NNPP model is able to perform path planning on novel maps. Experiments show that the guidance field generated by the NNPP model can significantly reduce the search time for optimal paths under the same hardware conditions, and the advantage of NNPP increases with the scale of the map.
keywords:
path planning, deep learning, planetary rovers, traversability cost, elevation map
[…]
[1,2,3,4,5,6]organization=State Key Laboratory of Robotics and Systems, addressline=Harbin Institute of Technology, city=Harbin, postcode=150001, state=Heilongjiang, country=China
1 INTRODUCTION
During planetary exploration missions, traversing rough terrain replete with obstacles poses challenges to rovers tasked with exploring diverse targets of interests. The surfaces of extraterrestrial bodies such as Mars and the Moon are beset with hazardous features including rocks, steep slopes and craters, while rovers are constrained by limited onboard energy reservessutoh2015right . Hence, identifying collision-free and energy-efficient path to destination sites emerges as a crucial requirement for planetary rovers aimed at extending their operational lifetimes and scientific output. Moreover, planetary rovers encounter challenging navigation issues during both short and long-range planning stages. The navigation problems are further exacerbated by the limited bandwidth and time lag for communication with Earth, which restricts the rovers to only one round-trip communication per sol (a Martian day that lasts approximately 24 hours and 40 minutes) in operational terms. Currently, the existing technology for driving Mars rovers mainly relies on images captured from the ground to plan movements. However, these plans can only be executed within the range visible to the rover. This underscores the importance of designing autonomous path planning algorithms for planetary rovers.
The planetary path planning problem can be classified into global path planning and local path planning. Global path planning requires determining the entire path from the rover’s starting position to its intended target using planetary surface images captured by orbiting satellites. On the other hand, local path planning is based on online maps generated by onboard sensors, making it suitable for unknown and dynamic environments. This paper focuses on addressing the global path planning problem for planetary rovers.
To address the navigation problem of mobile robots in rugged terrains such as planet surfaces, researchers initially proposed the idea of binary terrain classification into impassable obstacles and free traversal regionslaubach1998autonomous . In ettlin2006rough , Ettlin suggested combining binary motion planners with traversability classification by integrating rapidly-exploring random trees (RRT) and gradually increasing the threshold for binary terrain classification. The previous methods may be too conservative since classifying an obstacle as non-traversable can result in a suboptimal plan or even failure to find a path. Given the rough nature of the terrain, an obstacle may be traversable from a particular direction or at a specific velocity. The issue of environment representation is of paramount importance in the domain of rough terrain navigation for autonomous systems. Hence, an adequate model that captures the accessibility of the terrain is needed. In traditional binary approaches, the environment is binarized into traversable and non-traversable categories. However, in rough terrain navigation it is not clear how to binarize the environment in a meaningful and robust manner, given the continuum of terrain difficulties. Therefore, a continuous approach that assigns a continuous cost value for traversing different parts of the terrain should be considered. Such an approach would provide a more nuanced representation of the environment, thereby allowing for more informed navigation planning.
The accuracy of global path planning depends on the scene information contained within the map. This information includes the layout of the terrain, the presence of obstacles, environmental conditions, and other relevant factors. These elements are crucial for a successful plan as they contribute to the understanding of the environment, which is essential for determining an optimal route. Compared to RGB images obtained by orbiting satellites, using a stereo camera CCD to capture stereo image pairs in orbit and generating digital elevation model (DEM) data containing terrain height information can provide more accurate elevation data of the planetary surface. In uneven planetary terrain, map data that includes height information can better reflect the requirements of planetary rovers for terrain exploration.
Thus, this paper adopts DEM data as the map representation and formulates the global path planning problem as finding the optimal feasible path that avoids untraversable areas, including obstacles, and minimizes the risk of failure due to hazardous terrain morphology, as shown in Figure. 1.
The remainder of the paper is organized as follows. Section 2 introduces related work on path planning, focusing on learning-based path planning methods. Section 3 presents environment modeling approaches and cost map computation methods, where the path with the lowest total cost is identified as the optimal path. Section 4 elaborates on the proposed method, including the network structure, dataset preparation, and model evaluation metrics. Section 5 introduces the simulation experiments designed for the model, which demonstrates that the model can significantly reduce optimal path search time.
2 RELATED WORKS
Numerous algorithms have been proposed and refined to tackle the path planning problem, including the ${A^{\star}}$ algorithmhart1968formal , RRT-based algorithmslavalle2001randomized , artificial potential field (APF) based algorithmsniu2023virtual , and intelligent algorithms such as the genetic algorithmzhang2016dynamic , ant colony algorithmliu2017improved and reinforcement learning based algorithmsjaradat2011reinforcement . This section focuses on the mainstream methods of path planning and their applications in rough terrain navigation.
2.1 Heuristic method
Heuristic path planning algorithms utilize a heuristic function ${H\left(n\right)}$ to select the extended nodes, significantly reducing the total number of extended nodes and improving the efficiency of trajectory planningwang2021warehouse . In a Grid map environment, the ${A^{\star}}$ algorithm can effectively represent obstacles of arbitrary shapes in two-dimensional space. However, in a three-dimensional environment, the evaluation function becomes complicated with the increase in dimensionality. Thus, the ${A^{\star}}$ algorithm is less commonly used in high-dimensional space. To address this limitation, many studies have proposed optimization schemes based on ${A^{\star}}$ algorithmsastar2020 , such as ${D^{\star}}$, ${FieldD^{\star}}$, ${Theta^{\star}}$, Anytime Repairing ${A^{\star}}$ (${ARA^{\star}}$), and more.
In the elevation map scenario, Bo Zhangzhang2022path utilized the cumulative traversability cost between grids as the cost function for ${A^{\star}}$ algorithm, thereby enabling it to search for the shortest path in 3D uneven terrains.
2.2 Learning based methods
In recent years, modulated by the rapid development and widespread applications of artificial intelligence technologies, especially the advances within deep reinforcement learning (DRL), learning-based methods for path planning, obstacle detection, trafficability analysis and related path reasoning tasks have received increasing interest from researchers in both academia and industry hoping to enable more autonomous navigation for unmanned systems operating in complex environmentstai2017virtual zhou2019vision .
Based on the heuristic methods discussed in the previous section, many researchers have employed learning-based techniques to construct guiding functions that accelerate the convergence process of classical algorithms. This approach has proven to be highly effective. The ${NRRT^{\star}}$wang2020neural approach applies a convolutional neural network (CNN) model to guide the sampling-based RRT path search on a 2D grid map while satisfying different clearance and step size planning requirements. The same team of authors proposed the NEED modelwang2021robot , which predicts promising search regions to guide the search direction of planning algorithms. This approach significantly enhances algorithm performance on new problems while maintaining a prediction speed of over 100 Hz on a laptop for high-dynamics applications. Reinforcement learning (RL), compared to traditional control methods, has the notable advantages of not requiring environmental maps, possessing strong learning capabilities from data, and exhibiting high dynamic adaptability to changes in complex environmentsmoreira2020deep . Its deep neural network architecture can effectively process high-dimensional information from various sensors, while its reinforcement learning mechanism can perform continuous decision-making tasks in complex and dynamic environments. This makes it particularly suitable for local path planning in dynamic scenarios. Many researchers have attempted to apply reinforcement learning methods to local path planning in rough terrain. Q-learning has been applied to local path planning, and according to Reference zhang2022path , reinforcement learning methods can yield superior paths compared to traditional approaches. Reference josef2020deep indicates that compared to baseline methods, reinforcement learning techniques can improve the success rate of path planning. It can also be inferred that reinforcement learning methods consume more time compared to traditional approaches.
The time taken by the Q-learning method is almost double that of the APF method and approximately 15 times that of the conventional ${A^{\star}}$ method. The success rate of reinforcement learning (95%-96%) also impacts its applicability to global maps at a larger scale. For example, Reference zhang2022path reports that the proposed reinforcement learning method consumes nearly 11 times the time compared to traditional ${A^{\star}}$ methods when searching paths in a ${110\times 110}$ elevation map, while only improving path optimality by 20%. The proposed NNPP method in this paper not only greatly reduces the search time (shortened by 4.6 times) and algorithm optimality is nearly on par with ${A^{\star}}$ (reduced by 2%) compared to traditional ${A^{\star}}$ on the same hardware and a ${256\times 256}$ map, but NNPP’s performance advantage will further increase with the expansion of the map size.
3 ENVIRONMENTAL MODELING AND PROBLEM FORMULATION
Different environmental modeling approaches can have different effects on the optimal path search results. DEM is commonly used in many instances of path planning on uneven terrains, thus this chapter introduces DEM-based environmental modeling approaches as well as the problem formulation discussed in this paper.
3.1 Digital Representation of Elevation Map
In practical applications of CE4wang2020vision , digital elevation model (DEM) is used for map generation and path searching. The spatial distribution of an uneven terrain environment is characterized utilizing a group of three-dimensional vectors ${\left\{{x,y,z}\right\}}$ in which ${x}$ and ${y}$ represent the planar coordinates of mapping an environmental feature point within the uneven terrain while ${z}$ denotes the altitude of the point. Therefore, geomorphic attributes including aspect, gradient of slope, changes in elevation, and roughness can be adequately extracted via DEM.
3.2 Information Extraction of Uneven Terrain Environment
According to the traversability assessment method used by Yutu 2 lunar roveryan2019autonomous , geomorphic attributes of the ${i{\rm{-th}}}$ grid cell in DEM data are calculated by the Eq. 1. To calculate the roughness, a patch consisting of the ${i{\rm{-th}}}$ grid cell and its eight neighboring cells is used.
${T_{i}}={k_{s}}\cdot\frac{{{\varphi_{i}}}}{{{\varphi_{s}}}}+{k_{r}}\cdot\frac{{{r_{i}}}}{{{r_{s}}}}+{k_{f}}\cdot\frac{{\Delta h}}{{{H_{s}}}}$ (1)
The angle between the normal vector and horizontal plane of grid ${i}$ is defined as ${\varphi_{i}}$. The variable ${r}$ denotes the residual of the fitting distance from each DEM data point in a unit cell to the fitted plane in the corresponding patch, representing the roughness of the terrain region. The variable ${\Delta h}$ represents the height difference between each DEM data point in a unit cell and the fitted plane in the corresponding patch, representing the elevation difference in the terrain region. The traversability cost for each grid cell is calculated as shown in Figure. 9. In Eq. 1, ${\varphi_{s}}$ represents the maximum safe tilt angle that a rover can traverse through a single grid cell, ${H_{s}}$ represents the maximum obstacle height value, and rs represents the maximum roughness value that a rover can navigate through. The variables ${k_{s}}$, ${k_{r}}$, and ${k_{f}}$ are weighting coefficients, and their sum is equal to 1. Let the traversability cost for traveling between the two grids ${i}$ and ${j}$ be Eq. 2.
${T_{i,j}}=\sqrt{{{\left({{T_{i}}-{T_{j}}}\right)}^{2}}}$ (2)
We adopt the approach of modeling the environment by calculating the cost of passing through each pixel value because in rugged and uneven terrains, the shortest Euclidean path may be impassable, thus it is necessary to integrally describe the terrain environment by combining slope, roughness and other geometric features in order to find the optimal traversable path. After the aforementioned environmental modeling, the problem discussed in this paper can be abstracted as searching for a path from a given start point to end point in a 2.5D cost map such that the path has the minimum accumulated traversability cost.
4 NNPP ALGORITHM
4.1 Network Structure
In this section, we introduce a neural network model for path planning. The model takes as input a three-channel combination tensor, where the first channel represents a cost map as detailed in part B, with the grayscale value of each pixel indicating the cost of the robot passing through that pixel, with higher costs indicating pixels that are closer to being impassable. The second channel encodes the path planning starting point using a Gaussian representation, as described in part C. Similarly, the third channel represents the endpoint’s Gaussian encoding. The model’s output is a two-dimensional single-channel probability map, visualized as a color map in Figure. 2. Each pixel in the probability map stores a probability value ${p}$ between 0 and 1, indicating the probability that the pixel contains the optimal path. Higher probabilities imply the model is more inclined to believe the optimal path passes through that pixel.
The detailed structure of the model is shown in Figure. 2. It uses an encoder-decoder architecture comprising a backbone module, pyramid pooling module and decoder module. From a unique perspective, the role of the model is equivalent to performing semantic segmentation on the map containing the start and end points, segmenting out those pixels that represent feasible paths, while we also expect the model to have a very fast inference speed. Therefore, in designing the model, we drew inspiration from excellent examples of real-time semantic segmentation models.
The NNPP model employs the STDCfan2021rethinking as the backbone network. Within its multi-stage encoding architecture, the spatial resolution of feature maps is reduced by a factor of two after each STDC stage. As such, a high-level semantic description of the input map can be extracted as the final stage 5 output of the STDC backbone. The low-level features output by Stage3 and Stage4 will be fused with the high-level features in the decoder to enhance the model’s performance for path planning. Feature fusion across different semantic levels has been shown to be effective in improving performance for tasks requiring precise localization such as path planningwang2020neural , by incorporating both low-level spatial details and high-level semantic information. Therefore, in our model’s decoder, we fuse the low-level features from earlier stages with the high-level features, in order to leverage spatial details that can help produce more accurate path planning outputs.
To enhance global scene contextualization, PSPNetzhao2017pyramid incorporated a pyramid pooling module (PPM) that concatenates pooling representations at multiple scales prior to the convolution layers, thereby forming both local component features and global contextual representations. Compared to the original PPM, we simplify it to use three different pyramid scales, and remove the concatenate operation between the original feature map and the upsampled features, as in the original PPM. Our simplified PPM can reduce computational cost. Specifically, we reduce the number of pyramid scales from the original 4 to only 3. Previous work has shown that features at different pyramid levels become increasingly semanticfu2019dual . Therefore, a fewer number of pyramid levels, such as 3, can capture sufficient semantic information, while reducing computational cost.
Feature fusion modules are commonly employed in semantic segmentation networks to enhance feature representations. More specifically, fusing multi-level features is instrumental in achieving high segmentation accuracy. The proposed model draws inspiration from the design of pplitesegpeng2022pp , utilizing an Attention Fusion Module to amalgamate low-level features and high-level features from the decoder. The structure of the Attention Fusion Module (AFM)peng2022pp is shown in Figure. 10.
The decoder likewise comprises several stages, which are tasked with merging and upsampling features. At each stage, features of different levels are fused through the AFM module. As the problem can be viewed as a pixel-wise binary classification problem, a conventional convolution layer with only one output channels serves as the classifier, producing the corresponding probability of the pixel belonging to the foreground or background class.
4.2 Dataset Generation
In the field of semantic segmentation for planetary surfaces, generating annotated data and training labels can be a very time-consuming taskdai2022segmarsvit . In the context of path planning tasks on lunar or Martian surfaces, obtaining the true optimal path can also be challenging. To address this issue, NNPP does not require any manually labeled data. Instead, it uses the optimal path obtained by ${A^{\star}}$ as a label to train the model. The 20-meter resolution lunar DEM dataset from CE2TMap2015xin2020geometric was used to generate a ${256\times 256}$ map. The cost of traversing each pixel was calculated using the method outlined in Eq. 1.
The classical ${A^{\star}}$ algorithm is a graph search algorithm commonly utilized for global path planningxie20202 . It is capable of finding the shortest path for a robot to travel between two predetermined positions in a known static environment. To facilitate the rover’s automatic avoidance of untraversable areas caused by hazard geomorphic terrain, this paper defines the cost function of ${A^{\star}}$ to include the traversability cost as follows:
$F\left(n\right)=G\left(n\right)+H\left(n\right)$ (3)
$G\left(n\right)={L_{sn}}+\omega\cdot{T_{sn}}$ (4)
Where ${{L_{sn}}=\sum_{k=1}^{n-1}{l_{k,k+1}}}$ represents the accumulated Euclidean distance that from the starting grid ${s}$ to the current grid ${n}$, and ${{T_{sn}}=\sum_{k=1}^{n-1}{T_{k,k+1}}}$ is the accumulated traversability cost between the two grids. This paper set ${H\left(n\right)}$ to be the Euclidean heuristic function.
4.3 Detailed Training Process and Evaluation Metric
Positional encoding is a powerful technique for infusing data with information regarding its position via embedding. In natural language processing (NLP), this approach is often leveraged to ensure that a model is aware of the location of words within a sentence. Inspired by this technique, NNPP makes use of positional encoding to embed location information about the starting and ending points on a map. The resulting encoded output is concatenated with the map, generating a three-channel tensor that serves as the input for the model. After encoding the starting point using sine and cosine functions to represent its absolute positionvaswani2017attention , the model’s outputs were not feasible. However, by switching to relative position encoding, more desirable results were achieved. This is because in practical path planning, the system is less concerned with absolute positions and more focused on relative distances between points. Specifically, our interest lies in the relative position of each cell in the elevation map compared to the starting point and the goal point.
$g\left({x,y}\right)=\frac{1}{{2\pi{\sigma^{2}}}}{e^{-\frac{{{{\left({x-cx}\right)}^{2}}+{{\left({y-cy}\right)}^{2}}}}{{2{\sigma^{2}}}}}}$ (5)
In this paper, a two-dimensional Gaussian distribution function centered around the starting point is used as the position encoding, with each grid in the elevation map represented by its corresponding value in the Gaussian function. This value reflects the relative position of the grid with respect to the starting point. The same approach is taken for the ending point.
Eq. 5 represents the calculation method for the Gaussian distribution function value at ${(x,y)}$. With ${\sigma}$ set to 50, ${cx}$ represents the ${x}$-coordinate of either the starting or ending point, while ${cy}$ represents the ${y}$-coordinate. Figure. 3 depicts the visualization results of Gaussian position encoding with different ${\sigma}$ values. The point coordinates are set to ${(200,200)}$. When ${\sigma}$ is too small, the effect is unsatisfactory, while when ${\sigma}$ is too large, the model also cannot extract sufficient starting point information from the encoding results. Experiments in [subsequent chapters] demonstrate that ${\sigma}$ at ${1/5}$ of the map size yields satisfactory results.
In the field of path planning, scholars have proposed many evaluation methods for path qualityliu2023path , among which average path planning time is an important metric that reflects the search efficiency of the algorithm. In elevation maps, the proposed method uses the cumulative traverse cost of a path to measure its optimality, believing that an optimal path should have the minimum cumulative traverse cost. This paper employs a learning-based approach to predict regions within a given map that contain optimal paths, whereas existing evaluation metrics are designed for a complete path rather than a region. Consequently, there is a need for designing an evaluation metric that directly assesses the precision of the model predictions. Figure. 4 illustrates the results of segmenting the model outputs according to diverse thresholds ${td}$, with the black areas representing regions assigned probabilities lower than ${td}$ by the model. The designation ${{}^{\prime}connected^{\prime}}$ denotes whether the segmented regions unite the starting point and end point. The present study decreases ${td}$ gradually from high to low values while detecting whether the two points are connected, thereby determining an optimal binary threshold and the corresponding region segmentation map. Subsequently, the total number of pixels that the model deems to contain optimal paths within the segmented regions is calculated, with the metric evaluated using the following Eq. 6. ${Z_{label}}$ represents the area of the labeled path, ${Z_{model}}$ denotes the area segmented using the adaptive threshold, and the areas are measured in number of pixels.
$ModelMetric=\frac{{{Z_{label}}}}{{{Z_{model}}}}$ (6)
In summary, there are three evaluation metrics: the average path planning time ${(AT)}$, the cumulative traverse cost of a path ${(CC)}$, and ${(MM)}$ as defined in Eq. 6.
5 SIMULATION EXPERIMENTS
In the present section, we delve into intricate parameter tuning and hardware environment settings, designing three experiments to demonstrate the efficacy of the proposed method. As detailed in 4.2, a training set of 100,000 samples of resolution ${256\times 256}$ was synthesized, 5000 samples were procured for elevation evaluation, and an additional 1000 samples were sequestered for performance metric validation in the present experiments. The training, elevation and validation sets were mutually exclusive from each other. The optimizer of choice for parameter optimization was AdamW with an initial learning rate of 0.001 and a weight decay of ${{10^{-4}}}$. After training for 150 epochs, the model was able to predict traversable areas between start and end points on 1000 terrain samples that it had not been exposed to previously, showcasing its generalization capabilities.
5.1 Comparative Experiments with Conventional Methods at Different Scales
The computationally intensive nature of utilizing algorithm ${A^{\star}}$ to search for optimal paths within elevation maps stems from the stipulation that ${A^{\star}}$ algorithm necessitates constructing a global search graph in advance of searching. The graph encodes the traversal costs for each pixel, incorporating the traversal costs between adjacent pixels.
Table 1 presents the time consumption for pathfinding using ${A^{\star}}$ within maps of varying dimensions. The path length indicated in the table represents the quantity of pixels encompassed within the path as determined by ${A^{\star}}$. To more reasonably reflect the proportion of graphing time within the total search time, the same scaling operations are applied to the start and end points when the map is scaled proportionally.
Inspection reveals that graph construction accounts for approximately 85% of the total time consumed by the algorithm. This entails an onerous outlay of computational resources necessitated by the construction and subsequent traversal of the exhaustive search graph. Indeed, the area subtended by optimal paths is markedly smaller than the total map area, indicating that a sizable proportion of the search graph constructed by ${A^{\star}}$ may be classified as unvaluable with respect to deriving optimal routes. In other words, the global search graph encompasses vast regions irrelevant to the final path, implying inefficiency in the algorithm’s graph generation procedure.
However, the proposed approach allows the neural network to focus on those regions that provide high value to subsequent path reconstruction algorithms such as ${RRT^{\star}}$ and ${A^{\star}}$. By reducing the spatial redundancy in the search space, the network can learn a more discriminative representation that facilitates more efficient replanning. Taking ${A^{\star}}$ as an example, constructing the search graph only within the regions output by NNPP, rather than a global search graph, as illustrated in Figure. 5, can save significant time. Moreover, as the map size increases, the advantages of the neural network approach become more pronounced. By leveraging the ability of the network to focus the search in the most promising areas, the proposed localized search graph yields increasingly superior speedup and scalability compared to exhaustive search-based algorithms as the problem size grows.
The model was trained on data of size ${256\times 256}$. When evaluating the model’s performance on ${512\times 512}$ maps, as illustrated in Figure. 6, the maps were first downsampled via average pooling. Denoted by s and g are the start and goal positions respectively. Subsequently, the method described in 4.3 was applied to obtain the mask of the regions of interest to the model. Finally, these regions were resized back to the original map size via interpolation. This procedure allowed the model trained on smaller maps to generalize well to larger maps, by focusing its search only on the relevant areas while ignoring the rest. The ability to downsample the input during inference and selectively pay attention to important regions enabled by the proposed method was critical for achieving scalability. The search graph was constructed only within the black regions in Figure. 4. Subsequently, the ${A^{\star}}$ method was applied on the search graph to compute the start-to-goal path. In comparison, an exhaustive search graph was built on the original map, and the same ${A^{\star}}$ method was utilized to search for the optimal path. By constraining the search space to only the salient areas identified by the proposed method, the localized search graph approach achieved significantly higher efficiency and scalability compared to the exhaustive search baseline. The ability to prune irrelevant regions of the map based on the predicted importance masks allowed the search complexity to grow sublinearly with the map size, in contrast to the quadratic complexity of exhaustive search. This demonstrates the effectiveness of the proposed approach in focusing the search in the most promising regions to guide the path planning algorithm.
As shown in the Figure. 6, the masked graph making module accepts two inputs: the model output post masking and the original map. In the original map, regions of interest as discerned by the model are retained and delineated in black while the remainder is rendered entirely white, with the white regions signifying impassable obstacles. In this manner, ${A^{\star}}$ search for the optimal path is constrained to a relatively smaller region. Provided that the specified region encompasses the optimum path, ${A^{\star}}$ will return a 100% consistent path matching the given label.
As Figure. 5 illustrates, the blue bars represent the computational time incurred by the NNPP method via the workflow in Figure. 6 to derive the optimal path whereas the orange bars denote the time consumed by conventional ${A^{\star}}$ search directly on the map. Evidently, as the map size increases, the neural network more decisively shortens the search time. Figure. 5b indicates the performance metrics measured using Eq. 6 and corresponding standard deviations for the NNPP model at diverse resolutions. In larger scenes, employing the upscaling via NNPP methodology to yield outputs compromises considerable original map information, resulting in the degraded model performance metrics exhibited in the figures. Yet, as Figure. 5a demonstrates, despite this information loss, the neural network approach can still remarkably shorten the time required to search for the optimal path. Despite sacrificing some fidelity, the neural network method can substantially reduce the computation time for finding the optimal path, as Figure. 5a illustrates.
As shown in Table 2, we tested the model’s path planning performance on a validation set consisting of 1000 maps, wherein ${A^{\star}NN}$ denotes the approach using the procedure of Figure. 6, utilizing the model output to guide ${A^{\star}}$ for path replanning.
Figure. 11 illustrates several scenarios of path planning on a ${256\times 256}$ map. It can be observed that the model’s output effectively avoids hazard terrain (where whiter pixels indicate higher traversal costs, indicating the increased steepness). The model provides relatively safe regions. For certain scenarios, the model presents two alternative paths, indicating its ability to handle semantic information of the global map. During actual planning, the model selects the path with a higher probability value. The corresponding data for the five scenarios are presented in Table 3.
The data presented in the table substantiates the claim that the method proposed in this paper significantly reduces the time consumed in optimal path planning, without substantial compromise on the path’s quality. Specifically, on a ${256\times 256}$ grid map, the NNPP (Neural Network Path Planning) method has achieved a 4.6-fold acceleration in the path searching time of the ${A^{\star}}$ algorithm, with a mere 2.9% decrease in path quality. As the map size increases, the advantages of the NNPP method are expected to become even more pronounced.
5.2 The Effects of Different Gaussian Positional Encoding
In contrast to other works in the field, we feed the start and end point information, encoded via positional encoding, into the model through separate channels. This approach effectively ensures that the model accords sufficient weight to the locational information of the start and end points during prediction. Literaturekulvicius2021one also inputs the start and end points via separate channels, but only with a single pixel each. While the approach can effectively predict paths in small grid maps up to ${20\times 20}$, it struggles in larger and more complex scenes such as the elevation maps discussed herein. The experiments depicted in Figure. 3 demonstrate that appropriate positional encoding of the start and end point parameters can considerably enhance the model’s delineation of traversable areas, whereas utilizing single pixels without positional encoding leads to unsatisfactory performance.
In Figure. 3, ${G}$ represents ${Gaussian}$ position encoding and ${E}$ represents ${Euclidean}$ position encoding. ${1/7}$ indicates that ${\sigma}$ is taken as ${1/7}$of the map size in Eq. 5. The Euclidean encoding method calculates the Euclidean distance from each grid cell to the start or end point as the relative position encoding, which is then normalized to generate the final 2D encoding. It can be seen that the model prediction performs the best when ${\sigma}$ is around ${1/4}$ of the map size. When ${\sigma}$ is very small, it is close to no position encoding, resulting in a significant decrease in model prediction performance.
5.3 Application to Dynamic Environment
For autonomous navigation robots, global path planning and local path planning target different scenarios, which are based on the global map from external sensors and local map from onboard sensors, respectively. However, there is no fundamental difference between them. The same algorithm with proper modifications can be applied to both scenarios. The NNPP method introduced in this paper essentially learns the mapping relationship between the terrain fluctuations within a certain area measured by traversability cost and the optimal path given start and end points. Therefore, whether it is a global map or a local map, the NNPP is capable of predicting optimal path regions from uneven terrains. To validate this idea, in this section, we employ the NNPP trained on a global lunar map dataset for path planning in local Martian maps. To further simulate the presence of obstacles in the local maps, several artificial noise points are introduced into the local Martian Digital Elevation Model (DEM), as illustrated in Figure. 8.
Furthermore, the predictive performance of the model in dynamic environments is also evaluated. The red pentagrams represent the destination points, while the green rectangles represent the starting points. Figure. 8 ${a)}$ and ${e)}$ validate the path planning for reaching different destination points under the same scene and starting point, while Figure. 8 ${b)-d)}$ and ${f)-h)}$ validate the path planning for the same starting and destination points in varying scenes.
6 CONCLUSION AND FUTUREWORK
This paper presents a neural network-based approach for path planning on uneven planetary surfaces, offering a solution to the demand for fast path planning in the autonomous navigation of planetary rovers. Drawing inspiration from real-time semantic segmentation models, the study adopts a channel-wise separation of maps, starting points, and destination points specifically for the path planning scenario. Furthermore, the investigation explores the impact of different positional encoding methods on the predictive performance of the model. The final simulation experiments demonstrate that the proposed method in this paper significantly reduces the computation time required for searching the optimal path in an elevation map compared to the traditional ${A^{\star}}$ method. Without the need for GPU acceleration, the model achieves an inference speed of 17 frames per second (FPS), enabling CPU-based real-time path planning in computationally constrained vehicular systems operating in dynamic scenarios. The inference time can be observed in Figure. 5.
Future work will focus on three aspects. Firstly, efforts will be directed towards enhancing the precision of the model’s output in identifying optimal regions. Secondly, the development of model-guided path reconstruction methods will be pursued. As the model output provides a region rather than a specific path, subsequent steps of the algorithm will need to combine traditional methods such as ${A^{\star}}$ and RRT to reconstruct a feasible path from the starting point to the destination. This leads to the third area of future work, which involves designing an end-to-end network architecture that directly outputs viable paths.